Find a Doctor (FaD) is a small startup that helps people find a physician that best meets their needs (location, insurance accepted, etc.). During a "slow" time for FaD, 7 staff members take calls from customers. On average, one call arrives every 5 minutes. On average, each staff member spends 20 minutes with each customer. How long (in minutes) does a customer spend on average waiting on hold before he or she can start speaking to a representative? (Choose the answer closest to your calculations) Group of answer choices A. 11.4 minutes B. 15.2 minutes C. 0.7 minutes D. 1.2 minutes.
Answer #1

Arrival rate (A) = 5 minutes or 60/5 = 12 per hour

Service rate (S) = 20 minutes or 60/20 = 3 per hour

Number of Servers (M) = 7

A/S = 12/3 = 4 and M = 7

Average number of people waiting (Lq) = 0.18

Waiting time = Lq/A hours = Lq/A*60 min = 0.18/12*60 = 0.9 min

Ans C